Vector A = 6.0 m and points 30° north of east. Vector B = 4.0 m and points 30° south of west. The resultant
vector A + B is given by?
I have tried this Question many times and have not been able to figure it out, Please help me
1
Expert's answer
2019-08-27T09:36:00-0400

For this task first, lets find X (East) and Y (North) components of each vector, and then sum them up.

"x_A=6*\cos30^\circ=5.19""y_A=6*\sin30^\circ=3"

"x_B=-4*\cos30^\circ=-3.46""y_B=-4*\sin30^\circ=-2"

So, the resultant vector will have coordinates:


"x=x_A+x_B=1.73""y=y_A+y_B=1"

The length of the vector:


"L=\sqrt{x^2+y^2}=2"

Direction:


"\varphi=tan^{-1}{\frac{y}{x}}=30^{\circ}"

Answer: resultant vector has the length of 2.0m and points 30 degrees north of east